OpenStudy (anonymous):

Graph using the slope and y-intercept 2x+5y=-25

OpenStudy (anonymous):

first bring the equation to standard form move 2x to the other side, what do you get?

OpenStudy (anonymous):

How can you solved it then i will learn on the steps

OpenStudy (anonymous):

I first move 2x to the other side: 5y = -2x -25 then I divide by 5: y = -2/5x - 5 now it's in standard form

OpenStudy (anonymous):

My slope is = 2/5

OpenStudy (anonymous):

the fraction is correct it should be negative though

OpenStudy (anonymous):

since you moved 2x to the other side, the sign changed

OpenStudy (anonymous):

Then my y=2x/5+5

OpenStudy (anonymous):

2x+5y=-25 5y = -25 -2x y= -5 -2/5 x

OpenStudy (anonymous):

My Major problem is the to graph it ,can you graph it using slope and y intercept

OpenStudy (anonymous):